DETECTION OF HIGHLY PREDISSOCIATIVE LEVELS OF CH B ^{2}Σ^{−} STATE WITH TWO-COLOR RES-ONANT FOUR-WAVE MIXING SPECTROSCOPY

\underline{YUAN-PERN LEE}, AWADHESH KUMAR, CHIH-CHANG SHIAO, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 30043.

We demonstrate an application of two-color resonant four-wave mixing (TC-RFWM) spectroscopy to detect highly predissociative levels of the B ^{2}Σ^{−} state of CH in a hostile environment of an oxyacetylene flame. The probe and the grating wavelengths are in resonance with the A ^{2}Δ - X ^{2}Π and B ^{2}Σ^{−} - X ^{2}Π transitions, respectively. We measured 42 previously unobserved rovibronic lines of the 0-0 band and additional 30 lines of the 1-0 band of the B ^{2}Σ^{−} - X 2Π system to access the rotational quantum number N’ up to 20 and 12 of =0 and 1 of the B ^{2}Σ^{−} state, respectively. Inclusion of measured additional line positions of the 0-0 and the 1-0 bands yields spectral parameters of the B ^{2}Σ^{−} state significantly improved over those obtained previously with only non-predissociative lines; one additional rotational constant L_{v} is determined. Although power saturation is significant even at lowest applicable laser energies, we estimated lifetimes with studies of power dependence of line widths. The lifetimes of the highest detectable level of B ^{2}Σ^{−} v’ = 0 and 1 are 3±1 ps (N’ = 20) and 8±2 ps (N’ = 11), respectively. We will also present preliminary data of TC-RFWM of CH in supersonic jet; saturation dip is readily observable even at low laser energies.